# Privacy-Preserving Retrieval-Augmented Generation with Differential Privacy

### v2

@@ 4.1 Methodology
We use two question-answering benchmarking datasets for RAG: Trivia [19] and Natural Question (NQ) [22]. Each dataset consists of a list of pairs of question and answer lists, i.e., every question can have multiple answers. By following the standard evaluations in RAG [5, 24, 17], we use the Wikipedia dataset as the external data source from which a retriever finds relevant documents. For each dataset, we use a subset of 100 questions to manage the computational overhead  [^4] We first filter out questions to less than 20 ground truth documents in the Wikipedia dataset. If a question relates with only a few documents, DP algorithms will likely fail since replacing a document would change the output a lot. Then, we split the remaining questions into 4 bins with 20–29, 30–39, 40–49 and 50–59 relevant documents and sample 25 questions from each bin..
@@
The retriever we use is the Dense Passage Retriever (DPR) [21] which is built on top of BERT [6]. It finds relevant documents that are close to the question in the embedding space produced by BERT. We compare the following generator LLMs: OPT (1.3B) [42], Llama 3.1 (8B) [8], and Pythia (1.4B) [3]. We additionally report the result of OPT (2.7B), Llama 3.2 (1B), and GPT2-XL [32] in Appendix B.1.
@@
We observe the results under multiple total privacy budgets, $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. More specifically, we sweep $\epsilon_{\mathrm{total}}=2$ to $40$ and set $\delta_{\mathrm{total}}=10^{-4}$. Furthermore, we consider different per-token privacy budgets for our private algorithms: $\epsilon_{\mathrm{token}}=1,2,5$ and $\delta_{\mathrm{token}}=10^{-5}$. We consider the number of voters $m$ of $10$, $20$, $30$, $40$, and $50$ for VoteRAG, and $30$, $40$, and $50$ for DPVoteRAG and DPSparseVoteRAG so as to ensure reasonable privacy-utility tradeoff and computational overhead. For DPSparseVoteRAG, we set the threshold $\tau$ to be half of the number of voters, i.e., $\tau=m/2$. When we use the LimitedDomain mechanism to privately choose the most frequent token, we set their parameter $\bar{k}$ to be the number of voters, where $\bar{k}$ is the limited size of the domain to which we add the Gumbel noise. For voting algorithms, each voter receives 1 relevant document, i.e., $k=1$. The utility evaluation metric is the match accuracy [26, 2, 34, 43] which measures if the prediction to a question contains any of its answers.
4.2 Results
@@
Table 1: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
We take a closer look at the effects of the hyperparameters in Table 1 with OPT (1.3B) on Trivia dataset under different total privacy budgets $\epsilon_{\mathrm{total}}$. We provide the detailed results, as in Table 1, with other LLMs in Appendix B.1.
@@
5 Related Work
@@
[41] proposed an empirical privacy-preserving algorithm for RAG through the synthetic data generation, while our work studies privacy-preserving RAG in the framework differential privacy, which protects the privacy of each individual document with the theoretical guarantee. Differential privacy has been studied in many other tasks in large language models too. Prompt tuning helps tailor the LLM to new tasks from a (private) test-domain dataset. [16] and [7] study the DP mechanism on two different prompt tuning frameworks: prompt optimization and offsite prompt tuning [35]. In-context learning adapts to different tasks by illustrating some examples in the context as the task description. DP in-context learning considers the situation when the examples are picked from any private set. [36] tackles this problem by generating synthetic examples with DP and [38] solves the DP test query by generating the answers, both in a sample-and-aggregate fashion. The differentially private pretraining and finetuning of LLMs has been studied to address the privacy concern in the training data and memory is a large bottleneck when naively deploying DP-SGD [1]. [25] focuses on the pretraining stage which introduces ghost clipping to make DP-SGD more memory efficient. [39] explores finetuning in the parameter-efficient framework LoRA [18]. Notice that DP voting plays a crucial role in these sample-and-aggregate algorithms, including ours. A basic approach is to apply the Laplacian or Gaussian mechanism [13]. [28, 29] proposed a data-dependent privacy analysis, which can be tighter when the majority vote has a large margin over other options. We integrate the LimitedDomain mechanism for our algorithm, which addresses challenges when the voting domain is large [9]; the large vocabulary size in token voting is our main bottleneck.
@@
Our algorithms generate the answers token by token, where each token needs a query to the private dataset and consumes some privacy budget. In this paper, we set up the privacy parameters before the start of the algorithm and have a pre-set maximum number of tokens to generate. However, the number of tokens to generate is different per question and is unknown before the algorithm starts – it is possible that the number of generated tokens is much smaller than the pre-set number but we still need to pay the full pre-defined privacy cost. A line of work [33, 15, 23, 37] tries to measure the privacy budget in fully adaptive composition where the budget consuming can interact with the data. Especially, [37] gives an analysis for this fully adaptive setting which matches the tightness of advanced composition. The idea of fully adaptive composition sounds a fit to our problem, which allows us to “pay as we go”, rather than predefining the $\epsilon_{\mathrm{total}}$ before the generation process. We found the analysis for fully adaptive setting is effective for large number of steps and small budget per step, while in our algorithm the number of generated tokens would not be very large and each token generation needs a relatively large budget to guarantee the utility. This mismatch makes us stick with the advanced composition.
6 Conclusion and Future Work
@@
Acknowledgments
TK, RW and KC would like to thank National Science Foundation NSF (CIF-2402817, CNS-1804829), SaTC-2241100, CCF-2217058, ARO-MURI (W911NF2110317), and ONR under N00014-24-1-2304 for research support.
Appendix A Proof of Theorem 1
Proof.
For any two neighboring datasets $D$ and $D^{\prime}$, the retrieved datasets are $D_{x}=R(x,D;mk)$ and $D_{x}^{\prime}=R(x,D^{\prime};mk)$ and we denote $L_{x}=(d_{1},\cdots,d_{mk})$ and $L_{x}^{\prime}=(d_{1}^{\prime},\cdots,d_{mk}^{\prime})$. We only need to show for any set of outcomes $S$, $\Pr_{A,\pi}[A(L_{x}^{\pi})\in S]\leq\exp(\epsilon)\Pr_{A,\pi}[A((L_{x}^{\prime})^{\pi})\in S]+\delta.$ First of all, $D_{x}$ and $D_{x}^{\prime}$ have at most one different document (without considering the order). Therefore we can define another list of documents $L^{\prime\prime}_{x}=(d_{1}^{\prime\prime},\cdots,d_{mk}^{\prime\prime})$, such that $L^{\prime\prime}_{x}$ is some ranking of $D_{x}\textquoteright$ ’ and it differs at most one position from $L$. Notice that $L^{\prime}_{x}$ and $L^{\prime\prime}_{x}$ are only different at orders. Therefore $(L_{x}^{\prime})^{\pi}$ and $(L_{x}^{\prime\prime})^{\pi}$ have same distributions and as a consequence $A((L_{x}^{\prime})^{\pi})$ and $A((L_{x}^{\prime\prime})^{\pi})$ have same distributions.
Appendix B Additional Experimental Results
B.1 Main Results across Models and Datasets
@@
For completeness, we further present the detailed results, in the same form of the one provided in Table 1, with OPT (1.3B and 2.7B), Pythia (1.4B), Llama 3.1 (8B), Llama 3.2 (1B), and GPT2-XL on Trivia and NQ datasets in Tables 212.
Table 2: Average match accuracy comparison of our algorithms on NQ dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 3: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 4: Average match accuracy comparison of our algorithms on NQ dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 5: Average match accuracy comparison of our algorithms on Trivia dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 6: Average match accuracy comparison of our algorithms on NQ dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 7: Average match accuracy comparison of our algorithms on Trivia dataset with Llama 3.1 (8B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 8: Average match accuracy comparison of our algorithms on NQ dataset with Llama 3.1 (8B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 9: Average match accuracy comparison of our algorithms on Trivia dataset with Llama 3.2 (1B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 10: Average match accuracy comparison of our algorithms on NQ dataset with Llama 3.2 (1B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 11: Average match accuracy comparison of our algorithms on Trivia dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
Table 12: Average match accuracy comparison of our algorithms on NQ dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.

### v3

@@ IV Experiment
  4. 4.
    How do our algorithm protect against empirical privacy attack?
@@ IV-A Methodology
We use two question-answering benchmarking datasets for RAG: Trivia [joshi_triviaqa_2017] and Natural Question (NQ) [kwiatkowski_natural_2019]. Each dataset consists of a list of pairs of question and answer lists, i.e., every question can have multiple answers. By following the standard evaluations in RAG [chen_reading_2017, lewis_retrieval-augmented_2020, hsia2024ragged], we use the Wikipedia dataset as the external data source from which a retriever finds relevant documents. For each dataset, we use a subset of 100 questions to manage the computational overhead  [^5] We first filter out questions to less than 20 ground truth documents in the Wikipedia dataset. If a question relates with only a few documents, DP algorithms will likely fail since replacing a document would change the output a lot. Then, we split the remaining questions into 4 bins with 20–29, 30–39, 40–49 and 50–59 relevant documents and sample 25 questions from each bin..
In addition, we experiment with a realistic privacy-sensitive application, where the external corpus contains inherently private information. Chatdoctor Questions [li2023chatdoctor] consist of QA interactions between patients and doctors in the healthcare domain. We sample 100 patient questions from the original dataset as our test set. The external dataset consists of the remaining QA pairs from the original ChatDoctor dataset, excluding the 100 patient questions used for testing. Here is a document example:
Patient’s description: My son fell from bed heads on, and didn t vomit or pass out. However, we put him to sleep as this was his sleep time. After two hours he woke up, and we felt he had a fever. We gave him brufen… ### Doctor’s answer: Hi, according to me, I think since the kid fell and did not have vomiting, indicates there is no concussion. The fever is incidental, which can occur after a fall. This is normal. Give the kid paracetamol, probably every 6 hours once. If even after two days fever does not subside, kindly visit your pediatrician…. Take care.
@@
The retriever we use is the Dense Passage Retriever (DPR) [karpukhin_dense_2020] which is built on top of BERT [devlin_bert_2019-1]. It finds relevant documents that are close to the question in the embedding space produced by BERT. We compare the following generator LLMs: OPT (1.3B) [zhang_opt_2022-1], Llama 3.1 (8B) [dubey2024llama], and Pythia (1.4B) [biderman_pythia_2023-1]. We additionally report the result of OPT (2.7B), Llama 3.2 (1B), and GPT2-XL [radford_language_nodate] in Appendix -A.
@@
We observe the results under multiple total privacy budgets, $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. More specifically, we sweep $\epsilon_{\mathrm{total}}=2$ to $40$ and set $\delta_{\mathrm{total}}=10^{-4}$. Furthermore, we consider different per-token privacy budgets for our private algorithms: $\epsilon_{\mathrm{token}}=1,2,5$ and $\delta_{\mathrm{token}}=10^{-5}$. We consider the number of voters $m$ of $10$, 2020, $30$, $40$, and $50$ for VoteRAG, and $30$, $40$, and $50$ for DPVoteRAG and DPSparseVoteRAG so as to ensure reasonable privacy-utility tradeoff and computational overhead. For DPSparseVoteRAG, we set the threshold $\tau$ to be half of the number of voters, i.e., $\tau=m/2$. When we use the LimitedDomain mechanism to privately choose the most frequent token, we set their parameter $\bar{k}$ to be the number of voters, where $\bar{k}$ is the limited size of the domain to which we add the Gumbel noise. For voting algorithms, each voter receives 1 relevant document, i.e., $k=1$. The utility evaluation metric is the match accuracy [mallen_when_2023, asai_self-rag_2023-1, schick_toolformer_2023, zhang_retrievalqa_2024-1] which measures if the prediction to a question contains any of its answers.
IV-B Main Results
@@
TABLE I: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
We take a closer look at the effects of the hyperparameters in Table I with OPT (1.3B) on Trivia dataset under different total privacy budgets $\epsilon_{\mathrm{total}}$. We provide the detailed results, as in Table I, with other LLMs in Appendix -A.
@@
IV-C Empirical Privacy Evaluation
Figure 4: TPR-FPR curve of S2MIA for VoteRAG and DPSparseVoteRAG ($\varepsilon$=10) when the base LLMs are OPT (1.3B), Pythia (1.4B) and Llama 3.1 (8B).
To assess the degree of privacy protection offered by our proposed method, we evaluate the vulnerability of both a non-private RAG system and our privacy-preserving RAG system on the privacy-sensitive ChatDoctor dataset using membership inference attacks (MIA). Given a target document $x$ and a system $f_{D}$, an MIA computes a score $s(x,f)$ that reflects the likelihood of $x\in D$. Without loss of generality, we assume higher scores indicate a greater probability of membership. By applying the attack to two sets of documents (an in-distribution set $D_{\text{in}}\subset D$ and an out-of-distribution set $D_{\text{out}}$ with no overlap with $D$), we can derive a TPR–FPR curve and compute its AUC.
We adopt the membership score design from S2MIA [li2025generating]. In the ChatDoctor dataset, each document corresponds to a patient–doctor conversation. For a target document $x$, we extract the patient’s query $x_{t}^{q}$ and measure the similarity between the response $x_{t}^{r}$ generated by the tested RAG system and the doctor’s ground-truth answer $x_{t}^{g}$ in $x$. The similarity is quantified using the average precision score defined in BLEU [papineni2002bleu], which serves directly as the membership score in S2MIA.
Figure 4 presents the TPR–FPR curves and the corresponding AUC values. Without any privacy protection (VoteRAG), the attack is highly effective, yielding AUC values well above the diagonal baseline (0.5). In contrast, when querying our privacy-preserving system DPSparseVoteRAG with $\varepsilon=10$, the attack performance collapses to the naive baseline (AUC $\approx 0.5$). This demonstrates that our method effectively mitigates empirical privacy attacks while maintaining strong utility on the QA task, as shown in Figure 3.
IV-D More Analysis of DPSparseVoteRAG and DPVoteRAG
The length of generation.
The design of DPSparseVoteRAG might allow longer generation than DPVoteRAG because of the tighter composition from SVT. We empirically validated this intuition. Figure 5 shows the numbers of tokens generated by DPVoteRAG and DPSparseVoteRAG. As we expect by the design of DPSparseVoteRAG, we see DPSparseVoteRAG generates much more tokens than DPVoteRAG. This implies the effectiveness of the sparse vector technique in DPSparseVoteRAG to smartly spend privacy budget enabling long enough token sequences.
Figure 5: Average numbers of generated tokens for each $\epsilon_{\mathrm{token}}$ and $\epsilon_{\mathrm{total}}$ with OPT (1.3B) on Trivia dataset. We fix $m=50$. We report the means and standard deviations over three runs. We see DPSparseVoteRAG generates much more tokens than DPVoteRAG.
Effects of number of ground truth relevant documents.
Figure 6: Average match accuracy for questions with different numbers of ground truth relevant documents across baseline algorithms and our algorithms with OPT (1.3B) on Trivia dataset. We report the means and standard deviations over three runs. We see questions with more relevant documents tend to be answered correctly by our algorithm.
Figure 6 shows the performances for different numbers of ground truth relevant documents. We see questions with more relevant documents tend to be answered correctly by our algorithm.
V Related Work
@@
zeng-etal-2024-good proposed an empirical privacy-preserving algorithm for RAG through the synthetic data generation, while our work studies privacy-preserving RAG in the framework differential privacy, which protects the privacy of each individual document with the theoretical guarantee. Differential privacy has been studied in many other tasks in large language models too. Prompt tuning helps tailor the LLM to new tasks from a (private) test-domain dataset. hong2024dpopt and duan2024flocks study the DP mechanism on two different prompt tuning frameworks: prompt optimization and offsite prompt tuning [sordoni2023joint]. In-context learning adapts to different tasks by illustrating some examples in the context as the task description. DP in-context learning considers the situation when the examples are picked from any private set. tang2024privacypreserving tackles this problem by generating synthetic examples with DP and wu2024privacypreserving solves the DP test query by generating the answers, both in a sample-and-aggregate fashion. amin2024private proposes the aggregation based method to generate synthetic texts with DP, which applies the similar SVT idea of our methodology to save the budget for some tokens. The differentially private pretraining and finetuning of LLMs has been studied to address the privacy concern in the training data and memory is a large bottleneck when naively deploying DP-SGD [abadi2016deep]. li2022large focuses on the pretraining stage which introduces ghost clipping to make DP-SGD more memory efficient. yu2022differentially explores finetuning in the parameter-efficient framework LoRA [hu2022lora]. Notice that DP voting plays a crucial role in these sample-and-aggregate algorithms, including ours. A basic approach is to apply the Laplacian or Gaussian mechanism [dwork_algorithmic_2014]. papernot_semi-supervised_2017-1, papernot_scalable_2018-2 proposed a data-dependent privacy analysis, which can be tighter when the majority vote has a large margin over other options. We integrate the LimitedDomain mechanism for our algorithm, which addresses challenges when the voting domain is large [durfee_practical_2019-1]; the large vocabulary size in token voting is our main bottleneck.
@@
Our algorithms generate the answers token by token, where each token needs a query to the private dataset and consumes some privacy budget. In this paper, we set up the privacy parameters before the start of the algorithm and have a pre-set maximum number of tokens to generate. However, the number of tokens to generate is different per question and is unknown before the algorithm starts – it is possible that the number of generated tokens is much smaller than the pre-set number but we still need to pay the full pre-defined privacy cost. A line of work [rogers2016privacy, feldman2021individual, lecuyer2021practical, whitehouse2023fully] tries to measure the privacy budget in fully adaptive composition where the budget consuming can interact with the data. Especially, whitehouse2023fully gives an analysis for this fully adaptive setting which matches the tightness of advanced composition. The idea of fully adaptive composition sounds a fit to our problem, which allows us to “pay as we go”, rather than predefining the $\epsilon_{\mathrm{total}}$ before the generation process. We found the analysis for fully adaptive setting is effective for large number of steps and small budget per step, while in our algorithm the number of generated tokens would not be very large and each token generation needs a relatively large budget to guarantee the utility. This mismatch makes us stick with the advanced composition.
VI Conclusion and Future Work
@@
LLM usage considerations
We primarily use LLMs to refine the grammar and clarity of our writing, while the core ideas and research progress are developed independently through our own study and investigation.
-A Additional Experimental Results
@@
For completeness, we further present the detailed results, in the same form of the one provided in Table I, with OPT (1.3B and 2.7B), Pythia (1.4B), Llama 3.1 (8B), Llama 3.2 (1B), and GPT2-XL on Trivia and NQ datasets in Tables IIXV.
TABLE II: Average match accuracy comparison of our algorithms on NQ dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE III: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE IV: Average match accuracy comparison of our algorithms on NQ dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE V: Average match accuracy comparison of our algorithms on Trivia dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE VI: Average match accuracy comparison of our algorithms on NQ dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE VII: Average match accuracy comparison of our algorithms on Trivia dataset with Llama 3.1 (8B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE VIII: Average match accuracy comparison of our algorithms on NQ dataset with Llama 3.1 (8B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE IX: Average match accuracy comparison of our algorithms on Trivia dataset with Llama 3.2 (1B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE X: Average match accuracy comparison of our algorithms on NQ dataset with Llama 3.2 (1B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE XI: Average match accuracy comparison of our algorithms on Trivia dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
TABLE XII: Average match accuracy comparison of our algorithms on NQ dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.